Log(36/25)^3+3log2/9-log2=2log16/125 -proof
Answer by greenestamps(13334)



Straightforward application of the rules of logarithms....
(1) log(a)+log(b) = log(ab)
(2) log(a)-log(b) = log(a/b)
(3) log(a^b) = b*log(a)

And you didn't show the equation correctly; more parentheses are needed. It should be

log((36/25)^3) + 3log(2/9) - log(2) = 2log(16/125)

You can do the work.